# DexBench: Benchmarking LLMs for Personalized Decision Making in Diabetes Management

### v1

@@ Abstract
We present DexBench, the first benchmark designed to evaluate large language model (LLM) performance across real-world decision-making tasks faced by individuals managing diabetes in their daily lives. Unlike prior health benchmarks that are either generic, clinician-facing or focused on clinical tasks (e.g., diagnosis, triage), DexBench introduces a comprehensive evaluation framework tailored to the unique challenges of prototyping patient-facing AI solutions in diabetes, glucose management, metabolic health and related domains. Our benchmark encompasses 7 distinct task categories, reflecting the breadth of real-world questions individuals with diabetes ask, including basic glucose interpretation, educational queries, behavioral associations, advanced decision making and long term planning. Towards this end, we compile a rich dataset comprising one month of time-series data encompassing glucose traces and metrics from continuous glucose monitors (CGMs) and behavioral logs (e.g., eating and activity patterns) from 15,000 individuals across three different diabetes populations (type 1, type 2, prediabetes/general health and wellness). Using this data, we generate a total of 360,600 personalized, contextual questions across the 7 tasks. We evaluate model performance on these tasks across 5 metrics: accuracy, groundedness, safety, clarity and actionability. Our analysis of 8 recent LLMs reveals substantial variability across tasks and metrics; no single model consistently outperforms others across all dimensions. By establishing this benchmark, we aim to advance the reliability, safety, effectiveness and practical utility of AI solutions in diabetes care.
@@ 1 Introduction
Figure 1: DexBench spans 7 real-world tasks capturing realistic user needs in diabetes management.
@@
Therefore, we introduce DexBench  [^1] Dex is a play on the simple sugar dextrose, and Dexcom., a comprehensive evaluation framework tailored to the unique requirements of prototyping in the diabetes, glucose management and metabolic health domains. DexBench is the first large-scale LLM benchmark designed to evaluate model performance on real-world, patient-facing diabetes management tasks. Our benchmark spans 7 distinct task categories (see Figure 1), designed to encompass the breadth of decision-making questions individuals with diabetes ask. These range from basic glucose interpretation (“What is my time in range today?”), and behavioral associations (“Why did this salad cause a glucose spike?”), to decision making and planning (“What workouts from this past month consistently lower my glucose levels?”). We compile a rich dataset of one month of time-series CGM and behavioral data from 15,000 individuals across three populations: type 1 diabetes, type 2 diabetes, and prediabetes/general health and wellness. Using this data, we generate 360,600 personalized, contextual questions across the 7 task categories. To evaluate model performance, we develop multi-dimensional evaluation criteria for each task, covering 5 important metrics: accuracy, groundedness, safety, clarity, and actionability. Finally, we evaluate a diverse set of LLMs and find that no model consistently outperforms across all tasks and metrics, highlighting the need for continued improvement in LLMs for diabetes management.
@@
We present the following contributions: (1) We develop DexBench, a novel benchmark to evaluate LLMs on patient-facing diabetes management tasks created from wearable device data from 15,000 users across 3 diabetes populations. We generate 360,600 personalized, contextual questions, covering 7 real-world diabetes management tasks. (2) We develop a multi-dimensional evaluation framework for each task crafted by domain experts based on 5 key metrics: accuracy, groundedness, safety, clarity, and actionability. (3) We present comprehensive evaluations of 8 open-source and proprietary LLM models of differing sizes, purposes, and model families using DexBench. By establishing this benchmark, we aim to advance the reliability, safety and effectiveness of LLMs in metabolic health and diabetes, ultimately driving meaningful improvements for those living with diabetes. While focused on diabetes management, this framework is extensible to other domains involving wearable devices and continuous monitoring, including preventative care, fitness optimization and the management of other chronic conditions, e.g., hypertension, obesity, and sleep disorders. This benchmark also provides a foundation for evaluating LLMs on contextual reasoning tasks using complex, longitudinal time-series data across broader health and wellness applications.
2 DexBench
Figure 2: DexBench overview.
@@ 2 DexBench
In this section we present DexBench, a benchmark for evaluating LLMs on diabetes management decision making tasks. Figure 2 shows an overview and we walk through each component next.
@@ 2.2 Data Curation
To ensure DexBench reflects the diverse needs of real-world diabetes populations, we curated data from 15,000 individuals evenly distributed across three cohorts: prediabetes/health and wellness (HW), type 1 diabetes (T1D), and type 2 diabetes (T2D) (Figure 2b). Each user contributed 30 consecutive days of CGM data in mg/dL, recorded at 5-minute intervals. Data was collected between January and June 2025 and paired with time-aligned behavioral logs of meals, exercise, sleep, and other glucose metrics, as well as daily activity summaries (e.g., step count, average heart rate). Depending on the task, data were segmented into 1-day, 7-day, or 30-day windows. The most complete data from the 30 days were used for the 1-day and 7-day windows, meaning the consecutive days with the richest self-reported behavioral data across categories. We also generated user data from GlucoSynth, which provides highly realistic, differentially-private synthetic glucose traces (Lamp et al., 2023). Additional details, including LLM input formatting, are in Appendix A.3.
@@ 2.4 Model Evaluation
DexBench includes an evaluation framework to measure model performance across the full task suite (Figure 2e). Any LLM can be benchmarked by generating answers to task questions, which are then graded by an LLM evaluator, followed by verification by an expert human reviewer. We use Gemini 2.5 Pro as the LLM grader, with temperature and top-$p$ set to 0 for deterministic scoring. The grader uses a structured prompt to assign a binary score (0 or 1) for five metrics: accuracy, groundedness, safety, clarity, and actionability (see Table 5b in the Appendix). Each metric is designed to capture a distinct quality of model output. Accuracy measures factual correctness and logical soundness, with special checks for diabetes-specific terms (e.g., correct reference to glucose “in range” as 70–180 mg/dL). Groundedness evaluates contextualization, personalization, and fidelity to user data. Safety requires that outputs avoid harmful suggestions, and any medical recommendations, diagnoses, or prognoses. Clarity measures conciseness and readability, requiring a Flesch–Kincaid Grade level <8<8 (Kincaid et al., 1975), consistent with FDA medical device guidance, which recommends that key information be written at no higher than an eighth-grade reading level (FDA, 2001). Actionability judges whether responses provide useful, practical guidance. To ensure realistic and meaningful evaluation, we also define task-specific criteria and explicitly include them in the model prompts during answer generation for fair evaluation (see Appendix A.5 for specifics).
@@ 3 Results & Analysis
In this section, we report the comprehensive performance of a diverse set of LLMs on DexBench. We discuss LLMs evaluated and experimental settings in Section 3.1, present model performance results aggregated across all users and tasks in Section 3.2 and discuss additional analyses, i.e., model latency, impact of data input modality, and impact of model thinking budget in Section 3.3. Additional evaluation including per cohort and task-specific performance is in Appendix A.7-A.9.
@@ 3.1 Candidate LLMs & Experimental Settings
Table 2: Models evaluated with DexBench. The model suite spans a range of sizes, licenses, families, providers, and intended purposes.
@@
To establish baseline performance in DexBench, we evaluate eight different LLMs, shown in Table 2: Gemini 2.5 Pro (Comanici et al., 2025), GPT-5 (OpenAI, 2025), Gemini 2.5 Flash (Comanici et al., 2025), GPT-5-mini (OpenAI, 2025), Deepseek R1 0528 (DeepSeek-AI et al., 2025), Qwen 3 30B A3B Instruct (Yang et al., 2025), Llama 3.1 8B Instruct (Grattafiori et al., 2024), and MedGemma 4B Instruct (Sellergren et al., 2025). These models were selected based on availability due to privacy and legal constraints with the underlying user data, and to capture diversity across size, licensing, model families, and intended purpose. For all experiments, we report performance across all users, cohorts, and age groups (if applicable). Results are reported as the percent of model-generated answers that have passed a particular metric, along with standard error of mean (SEM). The SEM is calculated under a Bernoulli model, where for $n$ trials with $x$ successes the sample proportion is $p=\tfrac{x}{n}$, and $\mathrm{SEM}=\sqrt{\frac{p(1-p)}{n}}.$ Additional details are available in Appendix A.6.
@@ 3.2 Aggregated Results Across All Tasks
Table 3: DexBench Aggregated performance across all tasks. Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@
Figure 3 presents a summary comparison of the model performances across metrics for the 8 LLMs evaluated with DexBench. Overall, models tended to have strong performance on safety and actionability but were weaker on accuracy, groundedness and especially clarity. These findings aligned with our expectations as most models are likely tuned to provide safe outputs but often struggle to provide accurate, domain-specific calculations and to return outputs grounded in real data without hallucinations (Xu et al., 2025). Moreover, many models had weak performance on clarity because they struggled to provide responses at the appropriate reading level. The GPT-5 models in particular showed weaker performance on clarity, suggesting that they may not have been sufficiently optimized to adapt to the requested reading styles. On the other hand, DeepSeek R1 0528 had the strongest performance for clarity, but weaker performance for other metrics. This indicates DeepSeek generated simpler, more concise and understandable outputs, though they were not as accurate, grounded, safe, or actionable. Overall, while models reliably produced safe outputs, they consistently struggled with accuracy, groundedness, and domain-specific calculations, underscoring the tradeoff between safety and factual utility in user-facing AI systems.
@@
Table 4: Task-specific challenges all models faced when tested on DexBench.
@@ 4 Related Work
DexBench is the first benchmark designed to evaluate LLM performance on real-world, patient-facing diabetes management tasks. It is built on a large and diverse cohort of 15,000 individuals spanning type 1 diabetes, type 2 diabetes, and prediabetes/health and wellness populations. Unlike previous benchmarks, which often involve small cohorts and clinician-centric tasks, DexBench emphasizes personalized, subjective decision-making and aims to support the development of AI tools that empower individuals in their daily self-management of diabetes.
@@ 5 Conclusion & Limitations
DexBench has the following limitations: First, the curated dataset lacks detailed cohort demographics (e.g., age) beyond diabetes type, is missing some relevant features like insulin and medications, and relies on wearable and self-logged data, which can be sparse and noisy. Also, while we curated 7 representative tasks, they do not capture the full breadth and complexity of diabetes management decision-making. Future work will explore expanding the dataset to incorporate other features, and extending the benchmark to support a wider range of health contexts and decision-making scenarios. Finally, while DexBench is focused on diabetes management, the framework is extensible to other domains involving wearable devices and continuous monitoring, and provides a foundation for evaluating LLMs on contextual reasoning tasks using complex, longitudinal time-series data.
In conclusion, we present DexBench, the first benchmark for evaluating LLMs on real-world decision-making tasks in diabetes management. Our evaluation of 8 diverse LLMs reveals that while models like GPT-5 and Gemini 2.5 Pro exhibit potential, none consistently outperform the rest across all 7 tasks and 5 metrics. Our analysis highlights opportunities for improvement, including in diabetes related mathematics and advanced contextual reasoning. We release DexBench publicly for extensible prototyping and to improve the suitability of LLMs for diabetes management.
@@ A.1 Ethics Statement
This work adheres to ethical standards in data collection, model evaluation, and benchmark design. All data used in DexBench were de-identified and obtained with appropriate consent and institutional approvals, ensuring participant privacy and compliance with relevant regulations (e.g., HIPAA). The benchmark is designed to evaluate AI systems in a patient-facing context, with a strong emphasis on safety, groundedness, and actionability to mitigate potential harms. We do not deploy or recommend clinical use of the evaluated models; instead, our goal is to promote responsible development and transparent assessment of AI tools in diabetes care. We acknowledge the limitations of current LLMs and advocate for continued research to ensure equitable, safe, and effective AI solutions for diverse populations.
@@ A.2 Benchmark Release
To foster collaboration and accelerate progress in AI and LLM development for diabetes management, we plan to release the extensible DexBench benchmark, including the general evaluation framework codebase, as well as our analysis results.
@@ A.4 Additional Question Generation Details
Table 5: DexBench Metrics: a) question generation metrics and b) model evaluation metrics.
@@ A.8.1 Task 1 (Glucose Math)
Table 6: DexBench performance for Task 1 (Glucose Math). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ 2.5 Flash
Task 1 (Glucose Math) involved 9 question templates used to generate unique questions per user, resulting in 135,000 answers per model, and 1,080,000 evaluations across all 8 models. Table 6 reports the scores across metrics for each model, along with an average across all metrics for Task 1 (Glucose Math). Results indicate that GPT-5 outperformed all models for each metric, with a 7.9% increase from the second strongest performance (Gemini 2.5 Pro and GPT-5-mini). GPT-5-mini also had strong performance, illustrating that the GPT-5 family tested on DexBench were strong in the diabetes-specific mathematics category. DexBench gives us the opportunity to dive deeper, specifically into model performance for each type of question, which tests diverse aspects of diabetes-related math and metrics.

### v2

# DM-Bench: Benchmarking LLMs for Personalized Decision Making in Diabetes Management

@@ Abstract
We present DM-Bench, the first benchmark designed to evaluate large language model (LLM) performance across real-world decision-making tasks faced by individuals managing diabetes in their daily lives. Unlike prior health benchmarks that are either generic, clinician-facing or focused on clinical tasks (e.g., diagnosis, triage), DM-Bench introduces a comprehensive evaluation framework tailored to the unique challenges of prototyping patient-facing AI solutions in diabetes, glucose management, metabolic health and related domains. Our benchmark encompasses 7 distinct task categories, reflecting the breadth of real-world questions individuals with diabetes ask, including basic glucose interpretation, educational queries, behavioral associations, advanced decision making and long term planning. Towards this end, we compile a rich dataset comprising one month of time-series data encompassing glucose traces and metrics from continuous glucose monitors (CGMs) and behavioral logs (e.g., eating and activity patterns) from 15,000 individuals across three different diabetes populations (type 1, type 2, prediabetes/general health and wellness). Using this data, we generate a total of 360,600 personalized, contextual questions across the 7 tasks. We evaluate model performance on these tasks across 5 metrics: accuracy, groundedness, safety, clarity and actionability. Our analysis of 8 recent LLMs reveals substantial variability across tasks and metrics; no single model consistently outperforms others across all dimensions. By establishing this benchmark, we aim to advance the reliability, safety, effectiveness and practical utility of AI solutions in diabetes care.
@@ 1 Introduction
Figure 1: DM-Bench spans 7 real-world tasks capturing realistic user needs in diabetes management.
@@
Therefore, we introduce DM-Bench  [^1] DM stands for Diabetes Management., a comprehensive evaluation framework tailored to the unique requirements of prototyping in the diabetes, glucose management and metabolic health domains. DM-Bench is the first large-scale LLM benchmark designed to evaluate model performance on real-world, patient-facing diabetes management tasks. Our benchmark spans 7 distinct task categories (see Figure 1), designed to encompass the breadth of decision-making questions individuals with diabetes ask. These range from basic glucose interpretation (“What is my time in range today?”), and behavioral associations (“Why did this salad cause a glucose spike?”), to decision making and planning (“What workouts from this past month consistently lower my glucose levels?”). We compile a rich dataset of one month of time-series CGM and behavioral data from 15,000 individuals across three populations: type 1 diabetes, type 2 diabetes, and prediabetes/general health and wellness. Using this data, we generate 360,600 personalized, contextual questions across the 7 task categories. To evaluate model performance, we develop multi-dimensional evaluation criteria for each task, covering 5 important metrics: accuracy, groundedness, safety, clarity, and actionability. Finally, we evaluate a diverse set of LLMs and find that no model consistently outperforms across all tasks and metrics, highlighting the need for continued improvement in LLMs for diabetes management.
@@
We present the following contributions: (1) We develop DM-Bench, a novel benchmark to evaluate LLMs on patient-facing diabetes management tasks created from wearable device data from 15,000 users across 3 diabetes populations. We generate 360,600 personalized, contextual questions, covering 7 real-world diabetes management tasks. (2) We develop a multi-dimensional evaluation framework for each task crafted by domain experts based on 5 key metrics: accuracy, groundedness, safety, clarity, and actionability. (3) We present comprehensive evaluations of 8 open-source and proprietary LLM models of differing sizes, purposes, and model families using DM-Bench. By establishing this benchmark, we aim to advance the reliability, safety and effectiveness of LLMs in metabolic health and diabetes, ultimately driving meaningful improvements for those living with diabetes. While focused on diabetes management, this framework is extensible to other domains involving wearable devices and continuous monitoring, including preventative care, fitness optimization and the management of other chronic conditions, e.g., hypertension, obesity, and sleep disorders. This benchmark also provides a foundation for evaluating LLMs on contextual reasoning tasks using complex, longitudinal time-series data across broader health and wellness applications.
2 DM-Bench
Figure 2: DM-Bench overview.
@@ 2 DM-Bench
In this section we present DM-Bench, a benchmark for evaluating LLMs on diabetes management decision making tasks. Figure 2 shows an overview and we walk through each component next.
@@ 2.2 Data Curation
To ensure DM-Bench reflects the diverse needs of real-world diabetes populations, we curated data from 15,000 individuals evenly distributed across three cohorts: prediabetes/health and wellness (HW), type 1 diabetes (T1D), and type 2 diabetes (T2D) (Figure 2b). Each user contributed 30 consecutive days of CGM data in mg/dL, recorded at 5-minute intervals. Data was collected between January and June 2025 and paired with time-aligned behavioral logs of meals, exercise, sleep, and other glucose metrics, as well as daily activity summaries (e.g., step count, average heart rate). Depending on the task, data were segmented into 1-day, 7-day, or 30-day windows. The most complete data from the 30 days were used for the 1-day and 7-day windows, meaning the consecutive days with the richest self-reported behavioral data across categories. We also generated user data from GlucoSynth, which provides highly realistic, differentially-private synthetic glucose traces (Lamp et al., 2023). Additional details, including LLM input formatting, are in Appendix A.3.
@@ 2.4 Model Evaluation
DM-Bench includes an evaluation framework to measure model performance across the full task suite (Figure 2e). Any LLM can be benchmarked by generating answers to task questions, which are then graded by an LLM evaluator, followed by verification by an expert human reviewer. We use Gemini 2.5 Pro as the LLM grader, with temperature and top-$p$ set to 0 for deterministic scoring. The grader uses a structured prompt to assign a binary score (0 or 1) for five metrics: accuracy, groundedness, safety, clarity, and actionability (see Table 5b in the Appendix). Each metric is designed to capture a distinct quality of model output. Accuracy measures factual correctness and logical soundness, with special checks for diabetes-specific terms (e.g., correct reference to glucose “in range” as 70–180 mg/dL). Groundedness evaluates contextualization, personalization, and fidelity to user data. Safety requires that outputs avoid harmful suggestions, and any medical recommendations, diagnoses, or prognoses. Clarity measures conciseness and readability, requiring a Flesch–Kincaid Grade level <8<8 (Kincaid et al., 1975), consistent with FDA medical device guidance, which recommends that key information be written at no higher than an eighth-grade reading level (FDA, 2001). Actionability judges whether responses provide useful, practical guidance. To ensure realistic and meaningful evaluation, we also define task-specific criteria and explicitly include them in the model prompts during answer generation for fair evaluation (see Appendix A.5 for specifics).
@@ 3 Results & Analysis
In this section, we report the comprehensive performance of a diverse set of LLMs on DM-Bench. We discuss LLMs evaluated and experimental settings in Section 3.1, present model performance results aggregated across all users and tasks in Section 3.2 and discuss additional analyses, i.e., model latency, impact of data input modality, and impact of model thinking budget in Section 3.3. Additional evaluation including per cohort and task-specific performance is in Appendix A.7-A.9.
@@ 3.1 Candidate LLMs & Experimental Settings
Table 2: Models evaluated with DM-Bench. The model suite spans a range of sizes, licenses, families, providers, and intended purposes.
@@
To establish baseline performance in DM-Bench, we evaluate eight different LLMs, shown in Table 2: Gemini 2.5 Pro (Comanici et al., 2025), GPT-5 (OpenAI, 2025), Gemini 2.5 Flash (Comanici et al., 2025), GPT-5-mini (OpenAI, 2025), Deepseek R1 0528 (DeepSeek-AI et al., 2025), Qwen 3 30B A3B Instruct (Yang et al., 2025), Llama 3.1 8B Instruct (Grattafiori et al., 2024), and MedGemma 4B Instruct (Sellergren et al., 2025). These models were selected based on availability due to privacy and legal constraints with the underlying user data, and to capture diversity across size, licensing, model families, and intended purpose. For all experiments, we report performance across all users, cohorts, and age groups (if applicable). Results are reported as the percent of model-generated answers that have passed a particular metric, along with standard error of mean (SEM). The SEM is calculated under a Bernoulli model, where for $n$ trials with $x$ successes the sample proportion is $p=\tfrac{x}{n}$, and $\mathrm{SEM}=\sqrt{\frac{p(1-p)}{n}}.$ Additional details are available in Appendix A.6.
@@ 3.2 Aggregated Results Across All Tasks
Table 3: DM-Bench Aggregated performance across all tasks. Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@
Figure 3 presents a summary comparison of the model performances across metrics for the 8 LLMs evaluated with DM-Bench. Overall, models tended to have strong performance on safety and actionability but were weaker on accuracy, groundedness and especially clarity. These findings aligned with our expectations as most models are likely tuned to provide safe outputs but often struggle to provide accurate, domain-specific calculations and to return outputs grounded in real data without hallucinations (Xu et al., 2025). Moreover, many models had weak performance on clarity because they struggled to provide responses at the appropriate reading level. The GPT-5 models in particular showed weaker performance on clarity, suggesting that they may not have been sufficiently optimized to adapt to the requested reading styles. On the other hand, DeepSeek R1 0528 had the strongest performance for clarity, but weaker performance for other metrics. This indicates DeepSeek generated simpler, more concise and understandable outputs, though they were not as accurate, grounded, safe, or actionable. Overall, while models reliably produced safe outputs, they consistently struggled with accuracy, groundedness, and domain-specific calculations, underscoring the tradeoff between safety and factual utility in user-facing AI systems.
@@
Table 4: Task-specific challenges all models faced when tested on DM-Bench.
@@ 4 Related Work
DM-Bench is the first benchmark designed to evaluate LLM performance on real-world, patient-facing diabetes management tasks. It is built on a large and diverse cohort of 15,000 individuals spanning type 1 diabetes, type 2 diabetes, and prediabetes/health and wellness populations. Unlike previous benchmarks, which often involve small cohorts and clinician-centric tasks, DM-Bench emphasizes personalized, subjective decision-making and aims to support the development of AI tools that empower individuals in their daily self-management of diabetes.
@@ 5 Conclusion & Limitations
DM-Bench has the following limitations: First, the curated dataset lacks detailed cohort demographics (e.g., age) beyond diabetes type, is missing some relevant features like insulin and medications, and relies on wearable and self-logged data, which can be sparse and noisy. Also, while we curated 7 representative tasks, they do not capture the full breadth and complexity of diabetes management decision-making. Future work will explore expanding the dataset to incorporate other features, and extending the benchmark to support a wider range of health contexts and decision-making scenarios. Finally, while DM-Bench is focused on diabetes management, the framework is extensible to other domains involving wearable devices and continuous monitoring, and provides a foundation for evaluating LLMs on contextual reasoning tasks using complex, longitudinal time-series data.
In conclusion, we present DM-Bench, the first benchmark for evaluating LLMs on real-world decision-making tasks in diabetes management. Our evaluation of 8 diverse LLMs reveals that while models like GPT-5 and Gemini 2.5 Pro exhibit potential, none consistently outperform the rest across all 7 tasks and 5 metrics. Our analysis highlights opportunities for improvement, including in diabetes related mathematics and advanced contextual reasoning. We release DM-Bench publicly for extensible prototyping and to improve the suitability of LLMs for diabetes management.
@@ A.1 Ethics Statement
This work adheres to ethical standards in data collection, model evaluation, and benchmark design. All data used in DM-Bench were de-identified and obtained with appropriate consent and institutional approvals, ensuring participant privacy and compliance with relevant regulations (e.g., HIPAA). The benchmark is designed to evaluate AI systems in a patient-facing context, with a strong emphasis on safety, groundedness, and actionability to mitigate potential harms. We do not deploy or recommend clinical use of the evaluated models; instead, our goal is to promote responsible development and transparent assessment of AI tools in diabetes care. We acknowledge the limitations of current LLMs and advocate for continued research to ensure equitable, safe, and effective AI solutions for diverse populations.
@@ A.2 Benchmark Release
To foster collaboration and accelerate progress in AI and LLM development for diabetes management, we plan to release the extensible DM-Bench benchmark, including the general evaluation framework codebase, as well as our analysis results.
@@ A.4 Additional Question Generation Details
Table 5: DM-Bench Metrics: a) question generation metrics and b) model evaluation metrics.
@@ A.8.1 Task 1 (Glucose Math)
Table 6: DM-Bench performance for Task 1 (Glucose Math). Each entry shows the percentage of answers that passed a given metric $\pm$ SEM. Bold values indicate highest scoring model per metric.
@@ 2.5 Flash
Task 1 (Glucose Math) involved 9 question templates used to generate unique questions per user, resulting in 135,000 answers per model, and 1,080,000 evaluations across all 8 models. Table 6 reports the scores across metrics for each model, along with an average across all metrics for Task 1 (Glucose Math). Results indicate that GPT-5 outperformed all models for each metric, with a 7.9% increase from the second strongest performance (Gemini 2.5 Pro and GPT-5-mini). GPT-5-mini also had strong performance, illustrating that the GPT-5 family tested on DM-Bench were strong in the diabetes-specific mathematics category. DM-Bench gives us the opportunity to dive deeper, specifically into model performance for each type of question, which tests diverse aspects of diabetes-related math and metrics.